# Physics-Driven Construction of Compact Primitive Gaussian Density Fitting Basis Sets

### v1

@@ 2.3 Estimating energy contribution of DFBS primitives
In a single atom, DFBS must contain solid harmonics with $L$ up to $2L_{\text{OBS}}$ to obtain exact values for all possible matrix elements of the Hamiltonian in OBS, and the above regularization methods do not change this fact. However, it is well-known that the importance of DF AOs with the highest angular momenta is relatively low, at least for energies. Manually-optimized DFBSs typically restrict the $L_{\text{DFBS}}$ to significantly below $2L_{\text{OBS}}$, since only DFBS with $L$ up to $L_{\text{OBS}}+L_{\text{occ}}$ are needed for exact representation of $(ov|ov)$ integrals needed for the MP2 energy. In practice MP2-targeted DFBS are used successfully in the context of higher-order methods, like coupled-cluster. In ad hoc DFBS generators like AutoAux, the angular momenta of DFBS are restricted heuristically, i.e. $L_{\text{max,DF}}=\text{max}(2L_{\text{occ}},L_{\text{OBS}}+L_{\text{inc}})$, where $L_{\text{inc}}$ is fixed to 1 or 2 depending on $Z$61 To avoid the heuristics as much as possible, we attempted to design a first-principles energy-based model for further pruning of DFBS.
@@
Another question is how to obtain the 2-RDM. Our original efforts to model the importance of DFBS focused on heuristic models of RDMs (such as the Fermi-Dirac distribution). Thus in practice we constructed an estimator that requires a model for orbital occupancies $n_{p}\equiv\gamma_{p}^{p}$ only, rather than the full 2-RDM. To model the 2-RDM in terms of orbital occupancies, we use the Cauchy-Schwarz inequalities for the 2-RDM,82 which for positive $\gamma^{pq}_{pq}$ becomes simply:
@@
2-body energy pruning of the candidate DF AO pool $\mathcal{R}^{L}$ produced by Algorithm 1 proceeds as follows. If ${}^{2}\tilde{\bar{E}}^{\text{dx}}_{L}$ from (Eq. 17) is less than $Z\tau$, where $\tau$ is the target threshold, then all DF AOs of this angular momentum are discarded, else the DF AOs are sorted according to the order of importance provided by $|^{2}\bar{\epsilon}^{\text{dx}}_{X}|$. The candidate DF AOs are added to the $L$-channel of pruned DFBS, $\mathcal{D}^{L}$, until the difference between ${}^{2}\tilde{\bar{E}}^{\text{dx}}_{L}$ evaluated with $\mathcal{R}^{L}$ and with $\mathcal{D}^{L}$ is smaller than $Z\tau$. For an atom with ground-state configuration including occupied subshells of angular momenta $\leq L_{\text{occ}}$ DF channels $L\leq 2L_{\text{occ}}$ are pruned with $\tau=\tau_{1}$, the rest are pruned with $\tau=\tau_{2}$. Using the same values of $\tau$ for $Z=1$ (hydrogen) as for other $Z$ produced larger errors than desired due to the overpruning of the higher $L$ DF channels; thus, separate threshold $\tau_{\text{H}}$ is used for pruning DF channels of all angular momenta for H. The values of $\tau_{1}$, $\tau_{2}$, and $\tau_{\text{H}}$ will be determined heuristically in Section 4.1.
@@ 2.4 Correlated model of orbital occupancies
Our initial instinct was to model orbital occupancies in Eq. 14 by Fermi-Dirac distribution (with heuristics for the temperature). Unfortunately such an approach did not prove fruitful. Using atomic mean-field ensemble orbitals often used as initial guess for atomic density matrix would not suffice, since this will fail to generate nonzero populations for orbitals important in dynamical electron correlation. Therefore we developed a simple model for correlated 1-RDM of a neutral atom in an ensemble state using spin-opposite scaled (SOS) first-order Møller-Plesset perturbation of a model mean-field ensemble density. The mean-field density in OBS is obtained by one-shot diagonalization of OBS Fock matrix $\mathbf{F}$ computed from the atomic ensemble density generated by populating subshells of a minimal basis using the atom’s reference ground-state electron configuration. Such definition for the OBS Fock matrix is utilized in the Libint library and the MPQC program under the name “superposition of atomic densities” (SOAD) for generating guess Fock matrices for SCF. SOAD can be viewed as a simplified version of the popular SAD method85, with differences described for completeness in appendix A.
@@
produces energies and OBS AO coefficients of non-self-consistent ensemble orbitals. The orbitals are assigned spin-free occupancies $0\leq{\bf n}^{(0)}\leq 2$ using the same spherically-averaged ensemble definition used to construct the SOAD ensemble density and $\mathbf{F}_{\mathrm{SOAD}}$. Columns in ${\bf C}$ corresponding to zero and nonzero ensemble occupancies are deemed unoccupied ($a,b$) and occupied ($i,j$), respectively. These orbitals and the corresponding orbital energies are used to compute opposite-spin ensemble Møller-Plesset first-order amplitudes:
$t_{ab}^{ij}=\frac{\sqrt{n_{i}^{(0)}n_{j}^{(0)}}}{2}\frac{(ia|jb)}{\varepsilon_{i}+\varepsilon_{j}-\varepsilon_{a}-\varepsilon_{b}}.$ (21)
The corresponding second-order correlation contributions ${\bf n}^{(2)}$ to orbital occupancies are obtained straghtforwardly:
@@ 2.5 The MADF Algorithm
For a given (input) OBS consisting of contracted SHG AOs on an atom with atomic number $Z$ MADF (1) generates a “complete” set of primitive DF SHG AOs necessary to represent all products of uncontracted OBS AOs using the method of Lehtola62, (2) regularizes it by fusing pairs of shells with overly close exponents as described in Section 2.2, and (3) prunes out DF AOs not important for description of correlated 2-body energy using the method described in Section 2.3. Complete description of the algorithm is given in Algorithm 2. Its 4 model parameters (with 5 unique values) summarized in Table 1.
@@
$\tau_{\text{H}}$ 2-body energy threshold for the hydrogen atom $10^{-6}$
$\tau_{1}$ 2-body energy threshold for $Z>1$ and $L\leq 2l^{occ}$ $10^{-6}$ or $10^{-8}$ [^1] Optimal values of $\tau_{1}$ are different for nonrelativistic and relativistic OBSs, see Sections 4.1.2 and 4.1.3
$\tau_{2}$ 2-body energy screening threshold for $Z>1$ and $L>2l^{occ}$ $10^{-5}$
@@
1:function MADF($Z,\mathcal{O},\mathcal{M},\tau_{1},\tau_{2},\tau_{\text{H}},\zeta$) $\triangleright$ $Z$: atomic number; $\mathcal{O}$: (contracted) OBS; $\mathcal{M}$: MBS; $\tau_{\{1,2,\text{H}\}}\geq 0$; $\zeta\geq 1$
@@
3:     $\mathcal{C}\leftarrow\textsc{RegularizedCompleteDFBS}(\mathcal{P},\zeta)$triangleright Algorithm 1
4:     $\mathbf{F}_{\mathrm{SOAD}}\leftarrow\textsc{SoadFock}(Z,\mathcal{O},\mathcal{M})$triangleright appendix A
5:     $\mathbf{F}_{\mathrm{SOAD}}\mathbf{C}=\mathbf{SC\epsilon}$triangleright Eq. 20
6:     ${\bf n}\leftarrow\textsc{CorrelatedOccupancies}(\mathbf{C},\mathcal{M})$triangleright Eqs. 21, 22 and 24
7:     $L_{\text{occ}}\leftarrow\textsc{MaxAM}(\mathcal{M})$triangleright maximum angular momentum of occupied subshells.
8:     $L_{\text{max}}\leftarrow\textsc{MaxAM}(\mathcal{R})$triangleright maximum angular momentum of candidate pool of shells
9:     $\mathcal{D}\leftarrow\{\}$triangleright Initialize the target list of DF shells
@@
12:         if $Z\neq 1$ then
13:              if $L\leq 2L_{\text{occ}}$ then
14:                  $\tau\leftarrow\tau_{1}$
15:              else
16:                  $\tau\leftarrow\tau_{2}$
17:              end if
18:         else
19:              $\tau\leftarrow\tau_{\text{H}}$
20:         end if
21:         $\mathcal{r}\leftarrow\textsc{SelectByAM}(\mathcal{R},L)$triangleright Shells in $\mathcal{R}$ with angular momentum $L$
22:         $\{E_{\mathcal{r}},\{\epsilon_{\mathcal{r}}\}\}\leftarrow\textsc{Eval2EdxL}(\mathcal{r})$triangleright Evaluate Eqs. 17 and 18 using $\mathcal{r}$ as DFBS
23:         if $E_{\mathcal{r}}\geq Z\tau$ then
24:              $\mathcal{r}\leftarrow\textsc{SortBy}(\mathcal{r},\{|\epsilon_{\mathcal{r}}|\})$triangleright Sort $\mathcal{r}$ in order of decreasing $|{}^{2}\tilde{\bar{\epsilon}}^{\text{dx}}_{X}|$
25:              $E_{\mathcal{d}}\leftarrow 0,i\leftarrow 0$
26:              while $E_{\mathcal{r}}-E_{\mathcal{d}}\geq Z\tau$ and $i<\textsc{Size}(r)$ do
27:                  $\mathcal{d}\leftarrow\textsc{Concat}(\mathcal{d},\mathcal{r}[i])$
28:                  $\{E_{\mathcal{d}},\_\}\leftarrow\textsc{Eval2EdxL}(\mathcal{d})$triangleright Evaluate Eq. 17 using $\mathcal{d}$ as DFBS
29:                  $i\leftarrow i+1$
30:              end while
31:              $\mathcal{D}\leftarrow\textsc{Concat}(\mathcal{D},\mathcal{d})$
32:         end if
33:     end for
34:     return $\mathcal{D}$
35:end function
@@ 3 Technical details
The MADF algorithm is implemented in a developmental version of MPQC86; generation of base DFBS by Lehtola’s algorithm62 was implemented in Libint.87, 88 For optimization of MADF parameters and its benchmark testing, we employed the nonrelativistic and relativistic Hartree-Fock (HF) and second-order Møller-Plesset (MP2) methods. DF errors were computed relative to the HF and MP2 energies obtained with exact (four-center) two-electron matrix elements throughout. The exact two-component decoupled one-body Hamiltonian89, 90, 91 (1eX2C, or simply X2C) with spin-orbit coupling was used for all relativistic computations without empirical scaling corrections. The core orbitals were frozen in nonrelativistic MP2 computations with cc-pVXZ and def2- OBSs.
@@
The MADF model parameters for use with nonrelativistic OBS were determined on training set TS1 composed of 38 closed-shell systems with light elements: AlF3, Ar, BF3, BH3, C2H2, C2H4, C2H6, CH3OH, CH4, Cl2, CO2, COH2, CS2, F2, H2O, HCl, HF, HNO2, N2, cis-N2H2, trans-N2H2, Ne, NH3, NOCl, O2, O3, OF2, P2, PF3, PH3, S2, SF2, SiH4, SiO2, and SO2. The performance of the recommended model parameters was assessed on the standard G2 set of molecules92, containing 118 closed- and 30 open-shell systems. For training and assessment computations cc-pVXZ38, cc-pwCVXZ40(X = D, T, Q) and def2-SVP, def2-XZVP (X = T, Q)44 were used. Since there are no cc-pwCVXZ basis sets for Li, Be, and Na atoms, systems containing these atoms were skipped from the assessment of these OBSs.
@@
The MADF model parameters for use with relativistic OBS were determined on training set TS2 composed of 24 closed-shell systems with heavy elements: AgH, AsH3, AuH, BiH3, CuH, GaH3, GeH4, HAt, HBr, Hi, InH3, Kr, PbH4, PoH2, PtC, RuC, SbH3, ScH, SeH2, SnH4, TeH2, TlH3, Xe and Rn. The performance of the recommended model parameters was assessed on 2 sets: (a) the Ln54 benchmark set composed of 54 rare-earth compounds93, and (b) the set of 60 diatomics containing d-block (transition metal) atoms from Ref. 94 (we will refer to this set as Tm60). Due to the challenging SCF orbital optimization landscape and to make the assessment as robust and automated, for each OBS we did assessment on the Ln54 and Tm60 systems for which the exact (four-center) X2C-HF SCF converged in 30 or fewer SCF iterations. The training and assessment was performed with the following OBSs designed for spin-orbit relativity: X2C-SVPall-2c52, X2C-XZVPPall-2c53 (X = T, Q) and all-electron Dyall basis sets dyall-aeYz54, 55, 56 (Y = 2, 3, 4).
@@
The performance of MADF was assessed also against the AutoAux DFBS generator61 as implemented in BSE’s command-line interface41 . For cc-pVXZ, cc-pwCVXZ, and def2- family of OBSs we also compared against the corresponding manually-optimized DFBS cc-pVXZ-RIFIT, cc-pwCVXZ-RIFIT (X = D, T, Q) and def2-SVP-RIFIT and def2-XZVP-RIFIT (X = T, Q) for the G2 set for comparison. All geometries, except for the G2 set, were optimized with the PBE095 and def2-TZVP basis set using the Psi4’s96 geometry optimization module. For elements with atomic number greater than 36, the effective core potentials (ECPs)97 available for def2-TZVP were used for geometry optimization.
@@
All matrix pseudo-inverses and pseudo-inverse square roots were computed with Löwdin’s orthogonalization procedure98.

### v2

@@ 2.3 Estimating energy contribution of DFBS primitives
In a single atom, DFBS must contain solid harmonics with $L$ up to $2L_{\text{OBS}}$ to obtain exact values for all possible matrix elements of the Hamiltonian in OBS, and the above regularization methods do not change this fact. However, it is well-known that the importance of DF AOs with the highest angular momenta is relatively low, at least for energies. Manually-optimized DFBSs typically restrict the $L_{\text{DFBS}}$ to significantly below $2L_{\text{OBS}}$, since only DFBS with $L$ up to $L_{\text{OBS}}+L_{\text{occ}}$ are needed for exact representation of $(ov|ov)$ integrals needed for the MP2 energy. In practice MP2-targeted DFBS are used successfully in the context of higher-order methods, like coupled-cluster. In ad hoc DFBS generators like AutoAux, the angular momenta of DFBS are restricted heuristically, i.e. $L_{\text{max,DF}}=\text{max}(2L_{\text{occ}},L_{\text{OBS}}+L_{\text{inc}})$, where $L_{\text{inc}}$ is fixed to 1 or 2 depending on ZZ.61 To avoid the heuristics as much as possible, we attempted to design a first-principles energy-based model for further pruning of DFBS.
@@
Another question is how to obtain the 2-RDM. Our original efforts to model the importance of DFBS focused on heuristic models of RDMs (such as the Fermi-Dirac distribution). Thus in practice we constructed an estimator that requires a model for orbital occupancies $n_{p}\equiv\gamma_{p}^{p}$ only, rather than the full 2-RDM. To model the 2-RDM in terms of orbital occupancies, we use the Cauchy-Schwarz inequalities for the 2-RDM,85 which for positive $\gamma^{pq}_{pq}$ becomes simply:
@@
2-body energy pruning of the candidate DF AO pool $\mathcal{R}^{L}$ produced by Algorithm 1 proceeds as follows. If $^{2}\tilde{\bar{E}}^{\text{dx}}_{L}$ from (Eq. 17) is less than $Z\tau$, where $\tau$ is the target threshold, then all DF AOs of this angular momentum are discarded, else the DF AOs are sorted according to the order of importance provided by $|^{2}\bar{\epsilon}^{\text{dx}}_{X}|$. The candidate DF AOs are added to the $L$-channel of pruned DFBS, $\mathcal{D}^{L}$, until the difference between $^{2}\tilde{\bar{E}}^{\text{dx}}_{L}$ evaluated with $\mathcal{R}^{L}$ and with $\mathcal{D}^{L}$ is smaller than $Z\tau$. For an atom with ground-state configuration including occupied subshells of angular momenta $\leq L_{\text{occ}}$ DF channels $L\leq 2L_{\text{occ}}$ are pruned with $\tau=\tau_{1}$, the rest are pruned with $\tau=\tau_{2}$. For atoms with $L_{\text{occ}}=0$, namely H, He, Li and Be, to avoid overpruning of the higher $L$ DF channels, we consider $L_{\text{occ}}$ to be 1. The values of $\tau_{1}$ and $\tau_{2}$will be determined heuristically in Section 4.1.
@@ 2.4 Correlated model of orbital occupancies
Our initial instinct was to model orbital occupancies in Eq. 14 by Fermi-Dirac distribution (with heuristics for the temperature). Unfortunately such an approach did not prove fruitful. Using atomic mean-field ensemble orbitals often used as initial guess for atomic density matrix would not suffice, since this will fail to generate nonzero populations for orbitals important in dynamical electron correlation. Therefore we developed a simple model for correlated 1-RDM of a neutral atom in an ensemble state using spin-opposite scaled (SOS) first-order Møller-Plesset perturbation of a model mean-field ensemble density. The mean-field density in OBS is obtained by one-shot diagonalization of OBS Fock matrix $\mathbf{F}$ computed from the atomic ensemble density generated by populating subshells of a minimal basis using the atom’s reference ground-state electron configuration. Such definition for the OBS Fock matrix is utilized in the Libint library and the MPQC program under the name “superposition of atomic densities” (SOAD) for generating guess Fock matrices for SCF. SOAD can be viewed as a simplified version of the popular SAD method89, with differences described for completeness in Appendix A.
@@
produces energies and OBS AO coefficients of non-self-consistent ensemble orbitals. In this work, we have used the spin-free exact two-component 1-body (core) Hamiltonian (sf-1eX2C)90, 91, 92 to construct the $\mathbf{F}_{\mathrm{SOAD}}$ matrix to account for scalar relativistic effects. In general the model orbital energies of the atomic subshells obtained from Eq. 20 do not match the canonical order of Aufbau principle (Madelung rule) or even the order produced by self-consistent field in the given OBS. Thus the uncorrelated occupation numbers for the OBS ensemble orbitals ${\bf n}^{(0)}$ are obtained by mapping the spherically-averaged MBS ensemble occupancies ${\bf n}_{\mathrm{SOAD}}$ (Appendix A) via:
${\bf n}^{(0)}=\mathrm{id}\left(\mathbf{C}^{\dagger}\mathbf{S}_{\mathrm{OBS,MBS}}\right){\bf n}_{\mathrm{SOAD}},$ (21)
where $\mathbf{S}_{\mathrm{OBS,MBS}}$ is the overlap of OBS and MBS AOs, and $\mathrm{id}$ maps the real-valued matrix argument to the nearest identity matrix. Columns in $\bf C$} corresponding to zero and nonzero ensemble occupancies are deemed unoccupied ($a,b$) and occupied ($i,j$), respectively. These orbitals and the corresponding orbital energies are used to compute opposite-spin ensemble Møller-Plesset first-order amplitudes:
$t_{ab}^{ij}=-\frac{\sqrt{n_{i}^{(0)}n_{j}^{(0)}}}{2}\frac{(ia|jb)}{\varepsilon_{a}+\varepsilon_{b}-\varepsilon_{i}-\varepsilon_{j}}(1-e^{-\kappa(\varepsilon_{a}+\varepsilon_{b}-\varepsilon_{i}-\varepsilon_{j})})^{2}.$ (22)
The last factor on the right-hand side is the $\kappa$-regularizer used by Lee and Head-Gordon93 in the context of orbital-optimized MP2 method. Whereas in Ref. 93 heuristic tuning suggested $\kappa=1.45\,E_{\mathrm{h}}^{-1}$ as best at avoiding the singularities and accounting for the inaccuracy of the MP1 amplitude model. Here we set $\kappa=3\,E_{\mathrm{h}}^{-1}$ as the purpose of the regularizer here is to only avoid potential singularities due to possible degeneracies between occupied and unoccupied subshells; the inadequacy of the MP1 model for describing correlations of electrons in the partially occupied subshells is already partially addressed by avoiding internal excitations within such subshells. The corresponding second-order correlation contributions ${\bf n}^{(2)}$ to orbital occupancies are obtained straightforwardly:
@@ 2.5 The MADF Algorithm
For a given (input) OBS consisting of contracted SHG AOs on an atom with atomic number $Z$ MADF (1) generates a “complete” set of primitive DF SHG AOs necessary to represent all products of uncontracted OBS AOs using the method of Lehtola62, (2) regularizes it by fusing pairs of shells with overly close exponents as described in Section 2.2, and (3) prunes out DF AOs not important for description of correlated 2-body energy using the method described in Section 2.3. The complete description of the algorithm is given in Algorithm 2. Its 3 model parameters (with 4 unique values) are summarized in Table 1.
@@
$\tau_{1}$ 2-body energy threshold for $L\leq 2l^{occ}$ $10^{-6}$ (nr), $10^{-7}$ (rel) a
$\tau_{2}$ 2-body energy screening threshold for $L>2l^{occ}$ $10^{-5}$
@@
  • a
    Optimal values of $\tau_{1}$ are different for nonrelativistic and relativistic OBSs, see Sections 4.1.1 and 4.1.2
@@
1:function MADF($Z,\mathcal{O},\mathcal{M},\tau_{1},\tau_{2},\zeta$) $\triangleright$ $Z$: atomic number; $\mathcal{O}$: (contracted) OBS; $\mathcal{M}$: MBS; $\tau_{\{1,2\}}\geq 0$; $\zeta\geq 1$
@@
3:  $\mathcal{C}\leftarrow\textsc{RegularizedCompleteDFBS}(\mathcal{P},\zeta)$ $\triangleright$ Algorithm 1
4:  $\mathbf{F}_{\mathrm{SOAD}}\leftarrow\textsc{SoadFock}(Z,\mathcal{O},\mathcal{M})$ $\triangleright$ Appendix A
5:  $\mathbf{C}\leftarrow\mathbf{F}_{\mathrm{SOAD}}\mathbf{C}=\mathbf{SC\epsilon}$ $\triangleright$ Eq. 20
6:  ${\bf n}\leftarrow\textsc{CorrelatedOccupancies}(\mathbf{C},\mathcal{M})$ $\triangleright$ Eqs. 21, 22, 23 and 25
7:  $L_{\text{occ}}\leftarrow\textsc{Max}(\textsc{MaxAM}(\mathcal{M}),1)$ $\triangleright$ maximum angular momentum of occupied subshells.
8:  $L_{\text{max}}\leftarrow\textsc{MaxAM}(\mathcal{R})$ $\triangleright$ maximum angular momentum of candidate pool of shells
9:  $\mathcal{D}\leftarrow\{\}$ $\triangleright$ Initialize the target list of DF shells
@@
12:   if $L\leq 2L_{\text{occ}}$ then
13:     $\tau\leftarrow\tau_{1}$
14:   else
15:     $\tau\leftarrow\tau_{2}$
16:   end if
17:   $\mathcal{r}\leftarrow\textsc{SelectByAM}(\mathcal{R},L)$ $\triangleright$ Shells in $\mathcal{R}$ with angular momentum $L$
18:   $\{E_{\mathcal{r}},\{\epsilon_{\mathcal{r}}\}\}\leftarrow\textsc{Eval2EdxL}(\mathcal{r})$ $\triangleright$ Evaluate Eqs. 17 and 18 using $\mathcal{r}$ as DFBS
19:   if $E_{\mathcal{r}}\geq Z\tau$ then
20:     $\mathcal{r}\leftarrow\textsc{SortBy}(\mathcal{r},\{|\epsilon_{\mathcal{r}}|\})$ $\triangleright$ Sort $\mathcal{r}$ in order of decreasing $|{}^{2}\tilde{\bar{\epsilon}}^{\text{dx}}_{X}|$
21:     $E_{\mathcal{d}}\leftarrow 0,i\leftarrow 0$
22:     while $E_{\mathcal{r}}-E_{\mathcal{d}}\geq Z\tau$ and $i<\textsc{Size}(r)$ do
23:      $\mathcal{d}\leftarrow\textsc{Concat}(\mathcal{d},\mathcal{r}[i])$
24:      $\{E_{\mathcal{d}},\_\}\leftarrow\textsc{Eval2EdxL}(\mathcal{d})$ $\triangleright$ Evaluate Eq. 17 using $\mathcal{d}$ as DFBS
25:      $i\leftarrow i+1$
26:     end while
27:     $\mathcal{D}\leftarrow\textsc{Concat}(\mathcal{D},\mathcal{d})$
28:   end if
29:  end for
30:  return $\mathcal{D}$
31:end function
@@ 3 Technical details
The MADF algorithm is implemented in a developmental version of MPQC94; generation of base DFBS by Lehtola’s algorithm62 was implemented in Libint.95, 96 For optimization of MADF parameters and its benchmark testing, we employed the nonrelativistic and relativistic Hartree-Fock (HF) and second-order Møller-Plesset (MP2) methods. DF errors were computed relative to the HF and MP2 energies obtained with exact (four-center) two-electron matrix elements throughout. The exact two-component decoupled one-body Hamiltonian90, 91, 92 (1eX2C, or simply X2C) with spin-orbit coupling was used for all relativistic computations without empirical scaling corrections. The core orbitals were frozen in nonrelativistic MP2 computations with cc-pVXZ and def2- OBSs.
@@
The MADF model parameters for use with nonrelativistic OBS were determined on training set TS1 composed of 38 closed-shell systems with light elements: AlF3, Ar, BF3, BH3, C2H2, C2H4, C2H6, CH3OH, CH4, Cl2, CO2, COH2, CS2, F2, H2O, HCl, HF, HNO2, N2, cis-N2H2, trans-N2H2, Ne, NH3, NOCl, O2, O3, OF2, P2, PF3, PH3, S2, SF2, SiH4, SiO2, and SO2. The performance of the recommended model parameters was assessed on the standard G2 set of molecules97, containing 118 closed- and 30 open-shell systems. For training and assessment computations cc-pVXZ38, cc-pwCVXZ40(X = D, T, Q) and def2-SVP, def2-XZVP (X = T, Q)44 were used. Since there are no cc-pwCVXZ basis sets for Li, Be, and Na atoms, systems containing these atoms were skipped from the assessment of these OBSs.
@@
The MADF model parameters for use with relativistic OBS were determined on training set TS2 composed of 24 closed-shell systems with heavy elements: AgH, AsH3, AuH, BiH3, CuH, GaH3, GeH4, HAt, HBr, Hi, InH3, Kr, PbH4, PoH2, PtC, RuC, SbH3, ScH, SeH2, SnH4, TeH2, TlH3, Xe and Rn. The performance of the recommended model parameters was assessed on 2 sets: (a) the Ln54 benchmark set composed of 54 rare-earth compounds98, and (b) the set of 60 diatomics containing d-block (transition metal) atoms from Ref. 99 (we will refer to this set as Tm60). Due to the challenging SCF orbital optimization landscape and to make the assessment as robust and automated, for each OBS we did assessment on the Ln54 and Tm60 systems for which the exact (four-center) X2C-HF SCF converged in 30 or fewer SCF iterations. The training and assessment was performed with the following OBSs designed for spin-orbit relativity: X2C-SVPall-2c52, X2C-XZVPPall-2c53 (X = T, Q) and all-electron Dyall basis sets dyall-aeYz54, 55, 56 (Y = 2, 3, 4).
@@
The performance of MADF was assessed also against the AutoAux DFBS generator61 as implemented in BSE’s command-line interface41 . For cc-pVXZ, cc-pwCVXZ, and def2- family of OBSs we also compared against the corresponding manually-optimized DFBS cc-pVXZ-RIFIT, cc-pwCVXZ-RIFIT (X = D, T, Q) and def2-SVP-RIFIT and def2-XZVP-RIFIT (X = T, Q) for the G2 set for comparison. acCD calculations were performed using OpenMolcas version 24.10100. Since this version of OpenMolcas does not have an implementation of open-shell MP2, for comparison of MADF and acCD only the closed-shell subset of G2 set was used. We also show acCD vs MADF comparison only for the cc-pVXZ family of basis sets. All geometries, except for the G2 set, were optimized with the PBE0101 and def2-TZVP basis set using the Psi4’s102 geometry optimization module. For elements with atomic number greater than 36, the effective core potentials (ECPs)103 available for def2-TZVP were used for geometry optimization.
@@
All matrix pseudo-inverses and pseudo-inverse square roots were computed with Löwdin’s orthogonalization procedure104.